# We’ll Fix it in Post: Improving Text-to-Video Generation with Neuro-Symbolic Feedback

### v2

@@ 4 Methodology
Figure 3: Flowchart of video editing pipeline. The entire NeuS-E pipeline is three steps. Initially, figure 3(a) illustrates the original video generated by Pika. This video is missing ‘a sensor detects an error’ and ‘the system shuts down.’ NeuS-E takes the key frame and finds what events are missing. It then proposes an edited instruction to the key frame as shown in figure 3(b). Finally, a new video is generated with the modified key frame and instruction. The process repeats until the NeuS-V score goes beyond a given threshold, and the final video is shown in figure 3(c).
Given a synthetic video $\mathcal{V}$ generated by T2V models $\mathcal{M_{\text{T2V}}}:T,I\to\mathcal{V}$, where $T$ represents the input text prompt, and $I$ is the optional input image, our objective is to refine $\mathcal{V}$ to improve its temporal and logical consistency with the intended semantics of $T$. To achieve this, we introduce NeuS-E, a novel neuro-symbolic framework for improving T2V generation through structured neuro-symbolic feedback. We enhance the temporal and logical coherence of generated videos through the following steps:
  • Step 1: We first decompose the text prompt into the TL specification and a set of atomic propositions and then construct the video automaton using the video and VLM.
  • Step 2: We identify weak propositions and their corresponding frames by analyzing the video automaton’s structure and verification results, focusing on misaligned temporal or semantic behaviors.
  • Step 3: Finally, we convert the feedback into keyframe editing instructions (if required) and run the video editing pipeline. The video is iteratively refined until it meets the predefined threshold.
@@
First, we generate the video with the prompt without a input image: $\mathcal{V}=\mathcal{M_{\text{T2V}}}(T,\text{None})$. Then, we decompose $T$ into the TL specification $\Phi$ and the set of propositions $\mathcal{P}$. Given $\Phi$ and $\mathcal{P}$, we construct the video automaton as described in Definition 1 and verify it according to Definition 2. Finally, we perform a comprehensive neuro-symbolic verification of the generated video. Further details are provided in Fig. 2.
@@
In our running example demonstration (see Fig. 2), the original generation fails to capture “person is standing” and “person is walking away”. In this section, we discuss how to identify these weak propositions. The specification of our running example is in the form of ($A$wedge $B$) $\mathsf{X}$ $C$ $\mathsf{X}$ $D$, where $A=\text{person is meditating}$, $B=\text{lake shore}$, $C=\text{person is standing}$, and $D=\text{person is walking away}$. We aim to identify which component — $A$, $B$, $C$, or $D$ — is most misaligned in a generated video. Intuitively, we assess the video against partial specifications, such as ($B$) $\mathsf{X}$ $C$ $\mathsf{X}$ $D$, ($A$) $\mathsf{X}$ $C$ $\mathsf{X}$ $D$, ($A$wedge $B$) $\mathsf{X}$ $D$, and ($A$wedge $B$) $\mathsf{X}$ $C$, where each proposition’s confidence score is systematically adjusted. The proposition whose confidence adjustment causes the greatest shift in satisfaction probability reveals the weakest component.
Given $\mathcal{A}_{\mathcal{V}}$ and $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$, we aim to identify the weak proposition that most significantly degrades $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$. To achieve this, we systematically adjust the confidence score of each proposition independently and evaluate the resultant variation in $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$. Specifically, we adjust the confidence score that results in an adjusted confidence score set $\tilde{\mathbb{C}}$ as follows:
~=$\begin{split}\tilde{\mathbb{C}}&=\{\tilde{c}_{j,n}:\tilde{c}_{j,n}=1.0\;|\;p_{j}=\tilde{p}_{i};\\ &\quad\ \ \;\tilde{c}_{j,n}=c_{j,n}\;|\;p_{j}\neq\tilde{p}_{i},\;\forall p_{j}\in\mathcal{P}\}.\end{split}$ (4)
We then construct an evaluating video automaton $\tilde{\mathcal{A}}_{\mathcal{V},i}=\xi(\mathcal{P},\tilde{\mathbb{C}})$ that incorporates $\tilde{p}_{i}$ to reflect these adjustments and compute the new satisfaction probability $\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},i}\models\Phi]=\Psi(\tilde{\mathcal{A}}_{\mathcal{V},i},\Phi).$ Subsequently, we determine the weakest proposition $p^{*}_{i}$:
$\begin{split}p^{*}_{i}&=\\ &\arg\max_{i\in|\mathcal{P}|}\{\delta_{i}:\delta_{i}=\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},i}\models\Phi]-\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]\},\end{split}$ (5)
where $\delta_{i}\in[0,1]$ is the difference between the adjusted satisfaction probability $\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},i}\models\Phi]$ and the original $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$. The weakest proposition $p^{*}_{i}$ is the one that maximizes this difference, indicating the most significant impact.
@@
Given the weakest proposition $p^{*}_{i}$, we now localize its influence across $\mathcal{F}_{n}\in\mathcal{V}$ with respect to $p^{*}_{i}$. In our running example, after we identify the weakest proposition (e.g., “person is standing”) from the original generation, we determine the most impacted frames by the proposition, specifically the middle frame in the Fig. 2.
For each frame $\mathcal{F}_{n}$, we construct the evaluating automaton $\tilde{\mathcal{A}}_{\mathcal{V},n}$ for the video segment from the initial frame upto $\mathcal{F}_{n}$. We construct a new confidence measure $\hat{c}_{i,n}=1.0$ for the weakest proposition $p^{*}_{i}$ at $\mathcal{F}_{n}$. For all other frames $\{\mathcal{F}_{m}:m\in[0,n-1]\}$ and propositions $p_{i}\in\mathcal{P}$ including $p^{*}_{i}$, we set $\hat{c}_{i,n}=c_{i,n}+\gamma$, where $\gamma$ introduces controlled noise to enhance numerical stability in the model-checking process. Subsequently, we define the set of per-frame satisfaction probabilities =$\mathbb{Z}=\{\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},n}\models\Phi]=\Psi(\tilde{\mathcal{A}}_{\mathcal{V},n},\Phi)\ \mid\mathcal{F}_{n}\in\mathcal{V}\}$. Finally, we obtain the most impacted frame $\mathcal{F}^{*}_{n}$ by $\tilde{p}_{i}$ as follows:
@@ 4.4 Video Refinement Using Neuro-Symbolic Feedback
Through the above-mentioned neuro-symbolic feedback, that includes the baseline satisfaction probability $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$, the weakest proposition $p^{*}_{i}$, and the most impacted frame $\mathcal{F}^{*}_{n}$, we now aim to refine the video iteratively to enhance the temporal and logical consistency of the generated video $\mathcal{V}$.
@@
The frame $\mathcal{F}^{*}_{n}$ is mapped to its corresponding index in $\mathcal{V}$. We trim $\mathcal{V}$ at this index, retaining the segment from the initial frame to $\mathcal{F}^{*}_{n}$ as follows: $\mathcal{V}_{\text{trimmed}}=\mathcal{V}[0:n^{*}]$, where $n^{*}$ is the index of $\mathcal{F}^{*}_{n}$. This isolates the portion of the video most affected by $p^{*}_{i}$, preparing it for targeted regeneration.
Keyframe Image Editing.
Using the textual feedback for $p^{*}_{i}$ and $\mathcal{F}^{*}_{n}$, we edit the keyframe image. In our running example, the proposition “person is standing” is missing. Therefore, we modify $\mathcal{F}^{*}_{n}$ to depict an image of a standing person. We use a prompt-based image generation model OmniGen to edit $\mathcal{F}^{*}_{n}$:
^$\hat{\mathcal{F}}^{*}_{n}=\text{Omnigen}(I,T_{\text{img}}),$
where $T_{\text{img}}$ is the image edit prompt, $I$ is the input image, and ^$\hat{\mathcal{F}}^{*}_{n}$ is the edited keyframe. This step is optional. If we don’t edit the keyframe, the edited keyframe is the last frame from $\mathcal{V}_{\text{trimmed}}$.
@@
Now, we use an LLM to generate a new prompt $T_{\text{new}}$ for the next video segment following $\mathcal{V}_{\text{trimmed}}$. Additionally, we provide the weak proposition $p^{*}_{i}$ along with the prompt $T_{\text{video}}$ as input to the LLM. Once the new video generation prompt is obtained, we generate the new video segment as follows:
$\mathcal{V}_{\text{new}}=\mathcal{M}_{\text{VLM}}(T_{\text{new}},\hat{\mathcal{F}}^{*}_{n}).$
We present all prompts in the Appendix.
@@
Finally, the new video segment $\mathcal{V}_{\text{new}}$ is merged with $\mathcal{V}_{\text{trimmed}}$ and goes through the neuro-symbolic feedback process to obtain an updated $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$, the weakest proposition and the most impacted frame. The process iterates—regenerating and merging—until $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$ surpasses a predefined threshold or the maximum iteration limit is reached. We present the NeuS-E algorithm in the Appendix.
@@ 5 Experimental Setup
Following NeuS-V, we use GPT-4o [66] to decompose prompts into constituent propositions and generate temporal logic specifications. InternVL2.5-8B [11] serves as the VLM for obtaining per-frame confidence scores during automata construction. For further details on this process, we refer readers to NeuS-V [72].
Once our method identifies the weakest proposition and the most impacted frame, we use GPT-4o to generate both the video continuation prompt and the keyframe editing prompt. The keyframe, along with its corresponding edit prompt, is then processed by OmniGen [81] (which supports prompt-based image editing) to refine specific frames, typically by adding missing subjects. Then, the video is regenerated using the continuation prompt and the modified keyframe. This process iterates until either (1) the NeuS-V score surpasses 0.7, or (2) the number of iterations matches the number of propositions – ensuring that each weak proposition undergoes at least one round of refinement. Finally, we use MoviePy [97] for video trimming and stitching to integrate the edited video clips into a final MP4.
@@
We evaluate our method on both closed-source and open-source T2V models to demonstrate its broad applicability. Specifically, we conduct experiments on Gen-3 [70] by RunwayML and Pika-2.2 [49] by PikaArt as representatives of closed-source models, and CogVideoX-5B [87] as an open-source model. While our approach is designed to be model-agnostic and can be applied to any T2V system, we select these models due to their widespread use in the community and their ability to accept image inputs. This capability is essential for our editing pipeline, which iteratively improves temporal consistency by leveraging intermediate frames for controlled and coherent regeneration.
@@
We use the NeuS-V [72] prompt suite, which contains temporally extended prompts that challenge T2V models. It spans four themes (Nature, Human & Animal Activities, Object Interactions, Driving Data) and three complexity levels (Basic, Intermediate, Advanced). We evaluate both original and edited videos using the NeuS-V score to measure temporal fidelity and VBench, a widely used metric for synthetic video quality.
To establish an unbiased baseline, we include human evaluation alongside our quantitative metrics. Annotators are shown both the original video and the NeuS-E edited video and asked whether the latter better aligns with the caption. To ensure a focus on temporal fidelity, annotators are explicitly instructed to disambiguate general video quality from text-to-video alignment. For additional details on the evaluation setup, please refer to our Appendix.
@@ 6 Results
    Does NeuS-E improve text-to-video alignment without sacrificing video quality on temporally extended datasets?
@@
    How does iterative refinement through neuro-symbolic feedback compare to a step-by-step prompting strategy where prompts are carefully broken down and generated sequentially?
@@ 6.1 Improving Text-to-Video Generation with Neuro-Symbolic Feedback
As outlined earlier, we benchmark state-of-the-art T2V models, including Gen-3 and Pika as closed-source models, and CogVideoX-5B as an open-source alternative.
@@
We evaluate the before-and-after performance on the NeuS-V prompt suite in Table 1, measuring improvements across themes and complexity levels. By structuring results this way, we gain deeper insights into where and how NeuS-E enhances temporal alignment. At a glance, NeuS-E consistently improves text-to-video alignment across all themes and complexities. Notably, the improvement is more pronounced for higher-complexity prompts, where T2V models struggle the most. This suggests that while baseline models perform poorly on complex temporal relationships, our refinement method helps retain high temporal fidelity even as prompt difficulty increases. Additionally, we observe that Pika-2.2 excels at following video editing instructions, achieving an over 40% improvement in NeuS-V scores—outperforming other models in leveraging keyframe-based refinements.
Prompts Gen-3 Pika-2.2 CogVideoX-5B [87]
@@
Table 1: NeuS-V Score Improvements. Comparison of original and edited NeuS-V scores across different themes and complexity levels for Gen3, Pika, and CogVideoX models.
@@
To further validate the effectiveness of our approach, we conduct a human evaluation study, where annotators compare the original and edited videos and assess whether the latter better aligns with the prompt. They rate each comparison on a five-point scale: Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree. As shown in Figure 4, a significant percentage of annotators agreed that NeuS-E improves video alignment, correlating well with our benchmark evaluation results. Notably, Pika-2.2 shows the largest improvement, with nearly half of the videos rated as better after refinement. Other models also exhibit improvements, and crucially, the number of videos rated as worse remains consistently lower across all models. A key observation is the high proportion of neutral ratings. This is often due to (1) videos that were already well-aligned, making edits unnecessary, and (2) T2V models occasionally failing to follow regeneration instructions despite multiple iterations. The latter further highlights an inherent limitation of current T2V models in generating specific complex instructions.
@@
6.2 Ablation on Refinement Strategy – Is Neuro-Symbolic Feedback Key?
One could argue that our method is simply re-prompting a T2V model until the desired output is achieved. However, existing T2V models struggle to generate temporally coherent videos, even with repeated prompting. To determine whether our approach provides a meaningful advantage, we compare it against an alternative refinement strategy. Beyond standard re-prompting, we introduce a step-by-step generation baseline. Here, we leverage our proposition decomposition algorithm but without neuro-symbolic feedback. Instead of refining weak segments, this approach extends a video iteratively, generating each event sequentially by using the last frame of one generation as the first frame for the next.
Our results (Table 2) show that while step-by-step generation does improve temporal coherence, it falls significantly short of what neuro-symbolic feedback achieves. The key advantage of NeuS-E lies in its ability to extract two critical insights: (1) identifying weak propositions in the video and (2) determining the optimal time for those propositions to occur to best satisfy the temporal logic (TL) specification. Formal verification adds rigor to our approach. Additionally, we observe that step-by-step generation leads to longer videos on average, suggesting that without targeted neuro-symbolic feedback, the model produces redundant content rather than precisely correcting temporal inconsistencies.

### v4

@@ 4 Methodology
Given a synthetic video $\mathcal{V}$ generated by T2V models $\mathcal{M_{\text{T2V}}}:T,I\to\mathcal{V}$, where $T$ represents the input text prompt, and $I$ is the optional input image, our objective is to refine $\mathcal{V}$ to improve its temporal and logical consistency with the intended semantics of $T$. We introduce NeuS-E, a neuro-symbolic framework that operates through the following iterative steps:
  • Step 1: Decompose & Represent. The prompt $T$ is decomposed into a TL specification, and a video automaton is constructed from $\mathcal{V}$ using a VLM.
  • Step 2: Identify Errors. By analyzing the video automaton, the framework identifies weak propositions and their corresponding frames, pinpointing temporal misalignments.
  • Step 3: Refine & Iterate. The identified errors are converted into keyframe editing instructions to guide a video editing pipeline. This process is repeated until the refined video meets a predefined coherence threshold.
@@
First, we generate the video with the prompt without a input image: $\mathcal{V}=\mathcal{M_{\text{T2V}}}(T,\text{None})$. Then, we decompose $T$ into the TL specification $\Phi$ and the set of propositions $\mathcal{P}$. Given $\Phi$ and $\mathcal{P}$, we construct the video automaton as described in Definition˜1 and verify it according to Definition˜2. Finally, we perform a comprehensive neuro-symbolic verification of the generated video. Further details are provided in Figure˜2.
@@
[…]
Our method creates a hypothetical scenario for each proposition $\tilde{p}_{i}\in\mathcal{P}$ where we assume it is perfectly satisfied. Intuitively, we assess the video against partial specifications, such as ($B$) $\mathsf{X}$ $C$ $\mathsf{X}$ $D$, (AA) $\mathsf{X}$ $C$ $\mathsf{X}$ $D$, (AA $\wedge$ $B$) $\mathsf{X}$ $D$, and ($A$ $\wedge$ $B$) $\mathsf{X}$ $C$, where each proposition’s confidence score is systematically adjusted to find the weakest proposition. We achieve this by generating an adjusted confidence score set $\tilde{\mathbb{C}}$ that forces the confidence of the proposition under evaluation, $\tilde{p}_{i}$, to 1.0, while all other scores remain unchanged:
$\tilde{\mathbb{C}}=\{\tilde{c}_{j,n}:\tilde{c}_{j,n}=1.0\text{ if }p_{j}=\tilde{p}_{i};\quad\tilde{c}_{j,n}=c_{j,n}\text{ if }p_{j}\neq\tilde{p}_{i},\;\forall p_{j}\in\mathcal{P}\}.$ (4)
[…]
$p^{*}_{i}=\arg\max_{i\in|\mathcal{P}|}\{\delta_{i}\ :\ \delta_{i}=\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},i}\models\Phi]-\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]\}.$ (5)
The proposition $p^{*}_{i}$ with the highest $\delta_{i}$ value is the component most responsible for the video’s failure to align with the prompt, allowing us to target it for refinement.
@@
Given the weakest proposition $p^{*}_{i}$, we now localize its influence across $\mathcal{F}_{n}\in\mathcal{V}$ with respect to $p^{*}_{i}$. In our running example, after we identify the weakest proposition (e.g., “person is standing”) from the original generation, we determine the most impacted frames by the proposition, specifically the middle frame in the Figure˜2.
For each frame $\mathcal{F}_{n}$, we construct the evaluating automaton $\tilde{\mathcal{A}}_{\mathcal{V},n}$ for the video segment from the initial frame upto $\mathcal{F}_{n}$. We construct a new confidence measure $\hat{c}_{i,n}=1.0$ for the weakest proposition $p^{*}_{i}$ at $\mathcal{F}_{n}$. For all other frames $\{\mathcal{F}_{m}:m\in[0,n-1]\}$ and propositions $p_{i}\in\mathcal{P}$ including $p^{*}_{i}$, we set $\hat{c}_{i,n}=c_{i,n}+\gamma$, where $\gamma$ introduces controlled noise to enhance numerical stability in the model-checking process. Subsequently, we define the set of per-frame satisfaction probabilities $\mathbb{Z}=\{\mathbb{P}[\tilde{\mathcal{A}}_{\mathcal{V},n}\models\Phi]=\Psi(\tilde{\mathcal{A}}_{\mathcal{V},n},\Phi)\ \mid\mathcal{F}_{n}\in\mathcal{V}\}$. Finally, we obtain the most impacted frame $\mathcal{F}^{*}_{n}$ by $\tilde{p}_{i}$ as follows:
@@ 4.4 Video Refinement Using Neuro-Symbolic Feedback
We refine the video iteratively to enhance the temporal consistency of the generated video $\mathcal{V}$ through the above-mentioned neuro-symbolic feedback, which includes the baseline satisfaction probability $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$, the weakest proposition $p^{*}_{i}$, and the most impacted frame $\mathcal{F}^{*}_{n}$,
@@
We index the frame $\mathcal{F}^{*}_{n}$ in video $\mathcal{V}$ and trim the video to this index to obtain a segment to $\mathcal{F}^{*}_{n}$ as: $\mathcal{V}_{\text{trimmed}}=\mathcal{V}[0:n^{*}]$, where $n^{*}$ is the index of $\mathcal{F}^{*}_{n}$. This isolates the portion of the video most affected by $p^{*}_{i}$, preparing it for targeted regeneration.
@@
We use an LLM to generate a new prompt $T_{\text{new}}$ to generate the next video segment after $\mathcal{V}_{\text{trimmed}}$ by providing the weak proposition $p^{*}_{i}$ along with the original prompt $T_{\text{video}}$. We present all prompts in the Appendix.
@@
First, we merge the new video segment $\mathcal{V}_{\text{new}}$ with $\mathcal{V}_{\text{trimmed}}$, and repeat the neuro-symbolic feedback process as above to obtain the weakest proposition and the most impacted frame. This process is iterated until $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$ surpasses a predefined threshold or the maximum iteration limit is reached. We present the algorithm in the Appendix.
@@ 5 Experimental Setup
Following NeuS-V, we use GPT-4o [openai2023gpt4] to decompose prompts into constituent propositions and generate temporal logic specifications. InternVL2.5-8B [chen2024expanding] serves as the VLM for obtaining per-frame confidence scores during automata construction. For further details on this process, we refer readers to NeuS-V [sp_2025_CVPR].
Once our method identifies the weakest proposition and the most impacted frame, we use GPT-4o to generate the video continuation prompt. Then, the video is regenerated using the continuation prompt. This process iterates until either (1) the NeuS-V score surpasses 0.7, or (2) the number of iterations matches the number of propositions—ensuring that each weak proposition undergoes at least one round of refinement. Finally, we use MoviePy [MoviePy] for video trimming and stitching to integrate the edited video clips into a final MP4 (see Figure˜1).
@@
We evaluate our method on both closed-source and open-source T2V models to demonstrate its broad applicability. Specifically, we conduct experiments on Gen-3 [runway2024gen3] by RunwayML and Pika-2.2 [pikaai2024] by PikaArt as representatives of closed-source models, and CogVideoX-5B [yang2024cogvideox] as an open-source model. While our approach is designed to be model-agnostic and can be applied to any T2V system, we select these models due to their widespread use in the community and their ability to accept image inputs.
@@
We use the NeuS-V[sp_2025_CVPR] prompt suite, which contains temporally extended prompts that challenge T2V models. It spans four themes (Nature, Human & Animal Activities, Object Interactions, Driving Data) and three complexity levels (Basic, Intermediate, Advanced). We evaluate both original and edited videos using the NeuS-V score to measure temporal fidelity, and VBench[huang2024vbench], a widely used metric for synthetic video quality.
@@ 6 Results
    Does NeuS-E boost alignment without sacrificing aesthetics on temporally extended prompts?
@@
    How does iterative refinement through neuro-symbolic feedback compare to a sequential generation, where complex prompts are carefully broken down and generated sequentially?
@@ 6.1 Improving Text-to-Video Generation with Neuro-Symbolic Feedback
As outlined earlier, we benchmark state-of-the-art T2V models, including Gen-3 and Pika as closed-source models, and CogVideoX-5B [yang2024cogvideox] as an open-source alternative.
@@
We evaluate the before-and-after performance on the NeuS-V prompt suite in Table˜1, measuring improvements across different themes and complexity levels. By structuring results this way, we gain deeper insights into where and how NeuS-E enhances temporal alignment. At a glance, NeuS-E consistently improves text-to-video alignment across all themes and complexities. Notably, the improvement is more pronounced for higher-complexity prompts, where regular generation with T2V models struggle the most. This suggests that while baseline models perform poorly on complex temporal relationships, our refinement method helps retain high temporal fidelity even as prompt difficulty increases. Additionally, we observe that Pika-2.2 excels at following video editing instructions, achieving an over 40% improvement in NeuS-V scores—outperforming other models in leveraging keyframe-based refinements.
Table 1: NeuS-V Score Improvements. Comparison of original and edited NeuS-V scores across different themes and complexity levels for Gen3, Pika, and CogVideoX.
Prompts Gen-3 Pika-2.2 CogVideoX-5B
@@
-1em
@@
To further validate the effectiveness of our approach, we conducted a fully-blind, randomized A/B human evaluation. For each prompt, annotators were shown two videos (original vs. edited) in shuffled order and asked to judge which one better aligned with the caption on a five-point scale: Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree. As shown in Figure˜3, our edits are preferred in 52% of trials, closely mirroring the trend observed in our benchmark evaluations. Notably, Pika-2.2 shows the largest improvement, with nearly half of its videos rated as better after refinement. Other models also exhibit consistent gains, and importantly, the number of videos rated as worse remains low across all models. A substantial portion of neutral responses typically arises from (1) videos that were already well-aligned, making edits unnecessary, or (2) cases where T2V models failed to reliably follow regeneration instructions despite multiple iterations. The latter further highlights an inherent limitation of current T2V models in generating specific complex instructions.
@@
External Benchmark Validation.
To further validate that our improvements are not circular with respect to the NeuS-V metric, we evaluate NeuS-E on T2VCompBench [sun2025t2vcompbench], an independent compositional benchmark. On average, NeuS-E yields a $11\%$ improvement across the seven T2VCompBench dimensions, driven primarily by substantial gains in Action and Interaction, which are most closely tied to correcting missing or misordered events. We include category-wise improvements in Table˜2.
Table 2: T2VCompBench Scores Before and After Editing. Category-wise improvement trends for Pika-2.2 under NeuS-E refinement.
T2VCompBench Category Original Edited
Consist-Attr 0.590 0.700 (+0.110)
Dynamic-Attr 0.055 0.155 (+0.100)
Spatial 0.510 0.600 (+0.090)
Motion 0.270 0.390 (+0.120)
Action 0.510 0.660 (+0.150)
Interaction 0.580 0.710 (+0.130)
Numeracy 0.240 0.320 (+0.080)
Avg 0.394 0.505 (+0.111)
-1em
6.2 On Refinement Strategy – Is Neuro-Symbolic Feedback Key?
One could argue that our method is simply re-prompting a T2V model until the desired output is achieved. Conceptually, the NeuS-E pipeline contains two separable components: (1) diagnostic localization, which identifies the weakest proposition in the temporal logic specification and determines the frame where the failure occurs, and (2) execution, which regenerates the video segment conditioned on that diagnosis. Prompt rewriting occurs only in the second step. The key contribution of NeuS-E lies in the first step—the ability to formally diagnose where and why a video violates the temporal specification. To isolate this effect, we compare this effect with a baseline that performs sequential generation without this diagnostic signal (step-by-step generation). Our results (Table˜3), this baseline provides only marginal improvements, indicating that the gains arise primarily from neuro-symbolic localization rather than repeated prompting alone.
The key advantage of NeuS-E lies in its ability to extract two critical insights: (1) identifying weak propositions in the video and (2) determining the optimal time for those propositions to occur to best satisfy the temporal logic (TL) specification. Formal verification adds rigor to our approach. Additionally, we observe that step-by-step generation leads to longer videos, suggesting that without targeted neuro-symbolic feedback, the model produces redundant content rather than precisely correcting temporal inconsistencies.
A key strength of generating the full video first is that the T2V model establishes a unified global context across the entire timeline. In contrast, step-by-step generation treats segments in isolation or as auto-regressive extensions, which frequently leads to semantic drift (e.g., an actor’s appearance changing or the background shifting between steps). Furthermore, in the sequential approach, a minor hallucination in step $t$ becomes the ground truth for step t+1t{+}1; these errors accumulate and compound over the sequence. By generating the full video first and then applying surgical edits, NeuS-E retains the “strong” segments that benefit from global context and only subjects the “weak” segments to targeted, simplified regeneration.